OpenStudy (anonymous):

if A,B,C are angles made by vector 3i-6j+2k with positive direction of the coordinate axis then find cosA,cosB,cosC

OpenStudy (anonymous):

Divide each one of the components by the module of the vector and you will get the cosine of each angle. Module=SQROOT(9+36+4)=7 Then cos(A)=3/7; cos(B)=-6/7; cos(C)=2/7 assuming A, B and C are the angles of the vector with axis x, y and z respectively

OpenStudy (anonymous):

thanks